OpenStudy (anonymous):

A fighter plane horizontally at an altitude of 1.5 Km with speed of 720 Km/h passes directly overhead an anticraft gun. At what angle from the vertical should the gun be fired from the shell with speed of 400m/s to hit the plane in shortest time?

OpenStudy (jfraser):

@ what angle should the shell be fired so that it reaches an altitude of 1500m? This is a fairly simple projectile question

OpenStudy (anonymous):

u see the plane is too moving

OpenStudy (anonymous):

@JFraser

OpenStudy (across):

I assume the shell is fired the moment the plane is immediately above the gun?

OpenStudy (ujjwal):

Speed of plane= 200m/s let time be t Therefore 200\(\times\) t=ucos\(\theta\)\(\times\) t or, 200\(\times\) t=400cos\(\theta\) \(\times\)t or, \(\theta\)=60 drgree

OpenStudy (ujjwal):

*degree

OpenStudy (anonymous):

but @ujjwal the answer in my book is 30 degree

OpenStudy (ujjwal):

that is equal to 30 degree to vertical.

OpenStudy (ujjwal):

I used angle with horizontal

OpenStudy (anonymous):

yes u r correct

OpenStudy (anonymous):

thxxx

OpenStudy (ujjwal):

WC

OpenStudy (across):